How many grams of P2O5 and H2O will be produced in the reaction?



1
Expert's answer
2021-10-19T12:47:02-0400

Let use the following example:

If 17.0g of phosphine is mixed with 16.0g of O2 and reaction occurs. How many grams of P2O5 and H2O will be produced in the reaction?

2PH3 + 4O2 → P2O5 + 3H2O

M(PH3) = 34.00 g/mol

n(PH3) "= \frac{17.0}{34.0} = 0.5 \;mol"

M(O2) = 32.00 g/mol

n(O2) "= \frac{16.0}{32.0} = 0.5 \;mol"

According to the reaction equation for one mole of phosphine we need two moles of O2. So, O2 is a limitting reactant.

According to the reaction:

n(P2O5) "= \frac{1}{4}n(O_2) = \frac{1}{4} \times 0.5 = 0.125 \;mol"

M(P2O5) = 142.0 g/mol

m(P2O5) "= 0.125 \times 142.0 = 17.75 \;g"

According to the reaction:

n(H2O) "= \frac{3}{4}n(O_2) = \frac{3}{4} \times 0.5 = 0.375 \;mol"

M(H2O) = 18 g/mol

m(H2O) "= 0.375 \times 18.0 = 6.75 \;g"